A heavier mass m_1 and a lighter mass m_2 are 17.5
Answer #1

given,

distance = 17.5 cm

force = 9.4 * 10^-9 N

m1 + m2 = 5.95 kg

force = G * m1 * m2 / r^2

9.4 * 10^-9 = 6.674 * 10^-11 * m1 * m2 / 0.175^2

m1 = 0.844 kg

m2 = 5.1 kg